# A bound-preserving and conservative enriched Galerkin method for elliptic problems

### v1

@@ 4.4 Worst-case criterion for positivity preservation
Corollary 1 implies that for each iterate $w_{h}\in V_{h}$ defined via (11a), we have $w^{+}_{h}\in\mathcal{G}$ if $\|w^{0}_{h}\|_{0,\infty,\Omega}<\frac{b-a}{2}$. This section underlines that this criterion can always easily be satisfied if either $\beta$ is large enough, $\gamma$ is large enough, or $h$ is small enough. To this end, we start observing that using the inverse inequality, we get
@@
which can be combined with Lemma 2 and (4.3) to obtain
$\|\tilde{w}^{0}_{h}\|_{0,\infty,\Omega}\leq\,C\frac{h^{\beta-2-d/2}}{\gamma}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\big{(}\max(|a|,|b|)^{2}|\Omega|+\|w_{h}^{0}\|_{0,\Omega}^{2}\big{)}\right]^{\frac{1}{2}}\,.$
@@ 5 Convergence order estimates
In this section, we prove optimal order error estimates for the present bound-preserving EG method. The order of convergence appears as a result of the over-penalization of the jump terms. More precisely, the stability bound (12) proven in Theorem 1 is an error estimate in its own right, providing optimal convergence for $u_{h}^{0}$ if ${\beta\geq 4}$, since in such a case $u^{0}_{h}\to 0$ is sufficiently fast. Thus, the approximation properties of the current EG method stem from those of the continuous finite element subspace, while the piecewise constant enrichment is responsible for the local mass conservation. A similar interpretation of the best approximation properties can be used to argue for the convergence of DG methods, which converge optimally since the jump terms in the solutions converge to zero sufficiently quickly (the continuous solution does not have discontinuities).
We start by stating the following estimate for $u_{h}^{0}$. It is important to notice that this result is, in fact, a rewriting of Theorem 1.
Corollary 1.
Under the assumptions of Theorem 1, Let $u_{h}^{0}$ be the piecewise constant part of a solution $u_{h}{}$ of (9). Then, the following error estimate holds
$\|[\![u^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq\,C\,\frac{h^{{2\beta-4}}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|\right]\,.$ (19)
Remark 3.
Thanks to the broken Poincaré inequality from Lemma 2, we also have that $\|u_{h}^{0}\|_{0,\Omega}\leq C\,h^{\beta-{3/2}}$.
@@
Theorem 2.
Let us assume that $u\in H^{2}(\Omega)$ solves (2), $\beta\geq 4$, and that the assumptions of Theorem 1 hold. Then, the following error estimate holds
@@
$V^{+}_{h}=\{v_{h}\in V_{h}\colon v_{h}=P^{u^{0}_{h}}_{h}w_{h}\text{ for some }w_{h}\in V_{h}\}.$ (20)
@@
Proof.
As in the proof of [BGPV24, Th. 3.5], we can derive that
@@
Using that $P^{u^{0}_{h}}_{h}v_{h}=v_{h}$ and $Q^{u^{0}_{h}}_{h}v_{h}=0$ for $v_{h}\in V_{h}^{+}$ we exploit Lemma 5 to deduce that
@@
which implies the result. ∎
To exploit the variational inequality proven above, we must choose an appropriate test function. To achieve this, we introduce the Lagrange interpolation operator $\mathcal{I}_{h}:C^{0}(\bar{\Omega})\to V_{h}^{1}$ defined in [EG21a, Chapter 11], and define the test function
$\tilde{v}_{h}:=P^{u^{0}_{h}}_{h}(\mathcal{I}_{h}u).$ (21)
@@
Next, considering $\tilde{v}_{h}$ defined in (21), using the ellipticity of the bilinear form $a_{h}(\cdot,\cdot)$, the Cauchy-Schwarz inequality, the local trace-inverse result (14), the fact that $\tilde{v}_{h}-u_{h}^{+}$ is continuous, and that $[\![\tilde{v}_{h}-u]\!]={[\![P^{u^{0}_{h}}_{h}(\mathcal{I}_{h}u)]\!]=}[\![u_{h}^{0}]\!]$ we get
$\|\tilde{v}_{h}-u_{h}^{+}\|_{a_{h}}^{2}$ $=a_{h}(\tilde{v}_{h}-u_{h}^{+},\tilde{v}_{h}-u_{h}^{+})$
@@
which, after using the triangle inequality, proves the result. ∎
To prove the error, it only remains to bound the difference $u-\tilde{v}_{h}=u-P_{h}^{u_{h}^{0}}(\mathcal{I}_{h}u)$. We remark that $u_{h}^{0}$ has already been bounded in Corollary 1, so it only remains to bound the difference $u-{[P_{h}^{u_{h}^{0}}(\mathcal{I}_{h}u)]^{1}}$. The following result states that bound.
Lemma 13.
@@
Adding over the elements in the mesh yields the first inequality. The second one follows using the inverse inequality. ∎
Gathering all these preliminary results, we can finally prove Theorem 2.
Proof of Theorem 2.
Classical approximation properties of $\mathcal{I}_{h}$ (see, e.g., [EG21a]) yield
@@
In addition, using Lemma 13, the approximation properties of $\mathcal{I}_{h}$ and Corollary 1 we arrive at
@@
Gathering the last inequalities and using that $\beta\geq 4$ proves the error estimate. ∎
@@ 6 Numerical examples
This section presents various numerical examples to validate our theoretical findings and discuss implementation aspects. All computations were done using the finite element library Netgen/NGSovle, see [Sch14], www.ngsolve.org.
@@ 6.1 An iterative algorithm
To solve the nonlinear problem (9), we follow a similar strategy as was presented in [BGPV24], i.e., we consider using a Richardson-like iterative approximation. However, since an iteration of the (linearized) fully coupled problem defined on $V_{h}$ might lead to very ill-conditioned system matrices, see Section 6.2 below, we define a nested iterative scheme that is motivated by the fix point iteration (11a) introduced in Section 3. Each step in the iterative algorithm includes solving equations (11b) and (11c). While (11c) is linear and can be solved with classical tools (e.g., a direct solver), we introduce another iterative (sub) scheme to solve the non-linear problem (11b).
@@
Given a starting guess $u_{h}^{0}\big{|}_{0}\in V_{h}^{0}$ we solve for all $m=0,1,...$ the problem for $u_{h}^{0}\big{|}_{m+1}$: Step 1: Solve the following iteration to find $u_{h}^{1}\big{|}_{m+1}$: Given the starting guess $u_{h}^{1}\big{|}_{m}^{0}=u_{h}^{1}\big{|}_{m}$ and a damping parameter $\omega\in(0,1]$, for each $n=0,...$ we find $u_{h}^{1}\big{|}^{n+1}_{m}\in V_{h}^{1}$ such that for all $v_{h}^{1}\in V_{h}^{1}$ there holds $a_{h}(u_{h}^{1}\big{|}^{n+1}_{m},v_{h}^{1})=$ $a_{h}(u_{h}^{1}\big{|}_{m}^{n},v_{h}^{1})$ $+\omega\Big{(}(f,v_{h}^{1})-a_{h}([P_{h}^{u_{h}^{0}\big{|}_{m}}u_{h}^{1}\big{|}_{m}^{n}]^{1},v_{h}^{1})-s_{h}([Q_{h}^{u_{h}^{0}\big{|}_{m}}u_{h}^{1}|^{n}_{m}]^{1},v_{h}^{1})\Big{)},$ until $\|u_{h}^{1}\big{|}^{n+1}_{m}-u_{h}^{1}\big{|}^{n}_{m}\|_{0}\leq tol_{n}$. Let $N$ denote the number of iterations needed to reach the tolerance $tol_{n}$. We then set $u_{h}^{1}\big{|}_{m+1}=u_{h}^{1}\big{|}^{N}_{m}$. Step 2: To find $u_{h}^{0}\big{|}_{m+1}$, solve the problem: $a_{h}(u_{h}^{0}\big{|}_{m+1},v_{h}^{0})=(f,v_{h}^{0})-a_{h}(u_{h}^{1}\big{|}_{m},v_{h}^{0}),\quad\forall v_{h}^{0}\in V_{h}^{0}.$ Terminate the (outer) iteration when $\|u_{h}^{0}\big{|}_{m+1}-u_{h}^{0}\big{|}_{m}\|_{0}\leq tol_{m}$.
@@ 6.2 The condition number of the linearized problems
Let $\mathbb{A}$ denote the finite element matrix associated to the bilinear form $a_{h}(\cdot,\cdot)$, that is, the matrix that needs to be inverted when solving a linearized problem of (9). In the discussion that follows, we can neglect the stabilization term $s_{h}(\cdot,\cdot)$, as it only appears on the right-hand side of the inner iterations in Step 1 of the scheme presented in the last section. Further let $a_{h}^{i}(\cdot,\cdot)$ with $i\in\{0,1\}$ denote the restriction of the bilinear form $a_{h}$ on $V_{h}^{i}$, and let $\mathbb{A}^{i}$ denote the corresponding matrix. In the following, we discuss the condition numbers we expect for each system matrix, considering varying choices of $\beta$. By construction, the matrix $\mathbb{A}^{1}$ is the standard finite element matrix of linear Lagrange finite elements and is independent of $\beta$. The condition $\kappa(\mathbb{A}^{1})$ is then known to scale like $\mathcal{O}(h^{-2})$ which follows with standard techniques, see e.g. [EG21b, Chap. 28].
@@
Using the broken Poincaré inequality from Lemma 2 and the inverse trace inequality (14) we obtain
@@
$h\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0}\lesssim\underline{u}_{0}^{T}\mathbb{J}\underline{u}_{0}\lesssim h^{-1}\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0}\,,$ (22)
and using that $\kappa(\mathbb{M})\sim\mathcal{O}(1)$ (see, e.g., [EG21b, Proposition 28.6]) we conclude that $\kappa(\mathbb{J})\sim\mathcal{O}(h^{-2})$ (and thus is independent of the value of $\beta$). This further implies
@@
from which we also conclude $\kappa(\mathbb{J})\sim\mathcal{O}\Big{(}(1+h^{-(\beta+1)})/(1+h^{-\beta+1})\Big{)}$ which is bounded by $Ch^{-2}$ independently of the choice of $\beta$.
@@
$h$ $\kappa(\mathbb{A})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{1})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{0})^{-1}$ $eoc$
$0.5$ $1.52\cdot 10^{-3}$ (--- -) $2.03\cdot 10^{-2}$ (--- -) $1.28\cdot 10^{-1}$ (--- -)
@@
$h$ $\kappa(\mathbb{A})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{1})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{0})^{-1}$ $eoc$
$0.5$ $6.23\cdot 10^{-4}$ (--- -) $2.03\cdot 10^{-2}$ (--- -) $1.06\cdot 10^{-1}$ (--- -)
@@
$h$ $\kappa(\mathbb{A})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{1})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{0})^{-1}$ $eoc$
$0.5$ $9.51\cdot 10^{-5}$ (--- -) $2.03\cdot 10^{-2}$ (--- -) $6.69\cdot 10^{-2}$ (--- -)

### v2

@@ 4.4 Worst-case criterion for positivity preservation
Lemma 3.1 implies that for each iterate $w_{h}\in V_{h}$ defined via (11a), we have $w^{+}_{h}\in\mathcal{G}$ if $\|w^{0}_{h}\|_{0,\infty,\Omega}<\frac{b-a}{2}$. This section underlines that this criterion can always easily be satisfied if either $\beta$ is large enough, $\gamma$ is large enough, or $h$ is small enough. To this end, we start observing that using the inverse inequality, we get
@@
which can be combined with Lemma 4.2 and (4.19) to obtain
$\|\tilde{w}^{0}_{h}\|_{0,\infty,\Omega}\leq\,C\,h^{-d/2}\|\tilde{w}^{0}_{h}\|_{0,\Omega}\leq\,C\,h^{-d/2}\left\{\sum_{F\in{\mathcal{F}_{h}}}\frac{1}{h_{F}}\|[\![w_{h}^{0}]\!]\|^{2}_{0,F}\right\}^{\frac{1}{2}}\leq\,C\frac{h^{-d/2}}{\sqrt{\epsilon+\mu h^{2}}}\|w_{h}^{0}\|_{{\mathcal{F}_{h}}}\\ \leq\,C\,\frac{\min\{\epsilon^{-1/2},\frac{\mu^{-1/2}}{h}\}}{\gamma}h^{\beta-2-d/2}\left[\tfrac{C_{P}^{2}}{\mu}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\big(\max(|a|,|b|)^{2}|\Omega|+\|w_{h}^{0}\|_{0,\Omega}^{2}\big)\right]^{\frac{1}{2}}.$
@@ 5 Convergence order estimates
In this section, we prove optimal order error estimates for the present bound-preserving EG method. The order of convergence appears as a result of the over-penalization of the jump terms. More precisely, the stability bound (12) proven in Theorem 4.1 is an error estimate in its own right, providing optimal convergence for $u_{h}^{0}$ if $\beta\geq 4$}, since in such a case $u^{0}_{h}\to 0$ is sufficiently fast. Thus, the approximation properties of the current EG method stem from those of the continuous finite element subspace, while the piecewise constant enrichment is responsible for the local mass conservation. A similar interpretation of the best approximation properties can be used to argue for the convergence of DG methods, which converge optimally since the jump terms in the solutions converge to zero sufficiently quickly (the continuous solution does not have discontinuities).
We start by stating the following estimate for $u_{h}^{0}$. It is important to notice that this result is, in fact, a rewriting of Theorem 4.1.
Corollary 5.1.
Under the assumptions of Theorem 4.1, Let $u_{h}^{0}$ be the piecewise constant part of a solution $u_{h}{}$ of (9). Then, the following error estimate holds
(20) $\|[\![u^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq\,C\,\frac{h^{{2\beta-4}}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|\right]\,.$
Remark 5.2.
Thanks to the broken Poincaré inequality from Lemma 4.2, we also have that $\|u_{h}^{0}\|_{0,\Omega}\leq C\,h^{\beta-{3/2}}$.
@@
Theorem 5.3.
Let us assume that $u\in H^{2}(\Omega)$ solves (2), $\beta\geq 4$, and that the assumptions of Theorem 4.1 hold. Then, the following error estimate holds
@@
(21) $V^{+}_{h}=\{v_{h}\in V_{h}\colon v_{h}=P^{u^{0}_{h}}_{h}w_{h}\text{ for some }w_{h}\in V_{h}\}.$
@@
Proof 5.5.
As in the proof of [BarrenecheaGPV23, Thm. 3.5], we can derive that
@@
Using that $P^{u^{0}_{h}}_{h}v_{h}=v_{h}$ and $Q^{u^{0}_{h}}_{h}v_{h}=0$ for $v_{h}\in V_{h}^{+}$ we exploit Lemma 4.8 to deduce that
@@
which implies the result.
To exploit the variational inequality proven above, we must choose an appropriate test function. To achieve this, we introduce the Lagrange interpolation operator $\mathcal{I}_{h}:C^{0}(\bar{\Omega})\to V_{h}^{1}$ defined in [EG21-I, Ch.r 11], and define the test function
(22) $\tilde{v}_{h}:=P^{u^{0}_{h}}_{h}(\mathcal{I}_{h}u).$
@@
Next, considering $\tilde{v}_{h}$ defined in (22), using the ellipticity of the bilinear form $a_{h}(\cdot,\cdot)$, the Cauchy-Schwarz inequality, the local trace-inverse result (15), the fact that $\tilde{v}_{h}-u_{h}^{+}$ is continuous, and that $\tilde{v}_{h}-u]\!]={[\![P^{u^{0}_{h}}_{h}(\mathcal{I}_{h}u)]\!]=}[\![u_{h}^{0}]\!]$ we get
$\|\tilde{v}_{h}-u_{h}^{+}$ $\|_{a_{h}}^{2}$
$=a_{h}(\tilde{v}_{h}-u_{h}^{+},\tilde{v}_{h}-u_{h}^{+})$
@@
which, after using the triangle inequality, proves the result.
To prove the error, it only remains to bound the difference $u-\tilde{v}_{h}=u-P_{h}^{u_{h}^{0}}(\mathcal{I}_{h}u)$. We remark that $u_{h}^{0}$ has already been bounded in Corollary 5.1, so it only remains to bound the difference $u-{[P_{h}^{u_{h}^{0}}(\mathcal{I}_{h}u)]^{1}}$. The following result states that bound.
Lemma 5.8.
@@
Adding over the elements in the mesh yields the first inequality. The second one follows using the inverse inequality.
Gathering all these preliminary results, we can finally prove Theorem 5.3.
Proof 5.10 (Proof of Theorem 5.3).
Classical approximation properties of $\mathcal{I}_{h}$ (see, e.g., [EG21-I]) yield
@@
In addition, using Lemma 5.8, the approximation properties of $\mathcal{I}_{h}$ and Corollary 5.1 we arrive at
@@
Gathering the last inequalities and using that $\beta\geq 4$ proves the error estimate.
@@ 6 Numerical examples
This section presents various numerical examples to validate our theoretical findings and discuss implementation aspects. All computations were done using the finite element library Netgen/NGSovle, see [Schoberl14], www.ngsolve.org.
@@ 6.1 An iterative algorithm
To solve the nonlinear problem (9), we follow a similar strategy as was presented in [BarrenecheaGPV23], i.e., we consider using a Richardson-like iterative approximation. However, since an iteration of the (linearized) fully coupled problem defined on $V_{h}$ might lead to very ill-conditioned system matrices, see Section 6.2 below, we define a nested iterative scheme that is motivated by the fixed point iteration (11a) introduced in Section 3. Each step in the iterative algorithm includes solving equations (11b) and (11c). While (11c) is linear and can be solved with classical tools (e.g., a direct solver), we introduce another iterative (sub) scheme to solve the non-linear problem (11b).
@@
Given a starting guess $u_{h}^{0}\big|_{0}\in V_{h}^{0}$ we solve for all $m=0,1,$... the problem for $u_{h}^{0}\big|_{m+1}$: Step 1: Solve the following iteration to find $u_{h}^{1}\big|_{m+1}$: Given the starting guess $u_{h}^{1}\big|_{m}^{0}=u_{h}^{1}\big|_{m}$ and a damping parameter $\omega\in(0,1]$, for each $n=0,$... we find $u_{h}^{1}\big|^{n+1}_{m}\in V_{h}^{1}$ such that for all $v_{h}^{1}\in V_{h}^{1}$ there holds $a_{h}(u_{h}^{1}\big|^{n+1}_{m},v_{h}^{1})=$ $a_{h}(u_{h}^{1}\big|_{m}^{n},v_{h}^{1})$ $+\omega\Big((f,v_{h}^{1})-a_{h}([P_{h}^{u_{h}^{0}\big|_{m}}u_{h}^{1}\big|_{m}^{n}]^{1},v_{h}^{1})-s_{h}([Q_{h}^{u_{h}^{0}\big|_{m}}u_{h}^{1}|^{n}_{m}]^{1},v_{h}^{1})\Big),$ until $\|u_{h}^{1}\big|^{n+1}_{m}-u_{h}^{1}\big|^{n}_{m}\|_{0}\leq tol_{n}$. Let $N$ denote the number of iterations needed to reach the tolerance $tol_{n}$. We then set $u_{h}^{1}\big|_{m+1}=u_{h}^{1}\big|^{N}_{m}$. Step 2: To find $u_{h}^{0}\big|_{m+1}$, solve the problem: $a_{h}(u_{h}^{0}\big|_{m+1},v_{h}^{0})=(f,v_{h}^{0})-a_{h}(u_{h}^{1}\big|_{m+1},v_{h}^{0}),\quad\forall v_{h}^{0}\in V_{h}^{0}.$ Terminate the (outer) iteration when $\|u_{h}^{0}\big|_{m+1}-u_{h}^{0}\big|_{m}\|_{0}\leq tol_{m}$.
@@ 6.2 The condition number of the linearized problems
Let $\mathbb{A}$ denote the finite element matrix associated to the bilinear form $a_{h}(\cdot,\cdot)$, that is, the matrix that needs to be inverted when solving a linearized problem of (9). In the discussion that follows, we can neglect the stabilization term $s_{h}(\cdot,\cdot)$, as it only appears on the right-hand side of the inner iterations in Step 1 of the scheme presented in the last section. Further let $a_{h}^{i}(\cdot,\cdot)$ with $i\in\{0,1\}$ denote the restriction of the bilinear form $a_{h}$ on $V_{h}^{i}$, and let $\mathbb{A}^{i}$ denote the corresponding matrix. In the following, we discuss the condition numbers we expect for each system matrix, considering varying choices of $\beta$. By construction, the matrix $\mathbb{A}^{1}$ is the standard finite element matrix of linear Lagrange finite elements and is independent of $\beta$. The condition $\kappa(\mathbb{A}^{1})$ is then known to scale like $\mathcal{O}(h^{-2})$ which follows with standard techniques, see e.g. [EG21-II, Ch. 28].
@@
Using the broken Poincaré inequality from Lemma 4.2 and the inverse trace inequality (15) we obtain
@@
(23) $h\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0}\lesssim\underline{u}_{0}^{T}\mathbb{J}\underline{u}_{0}\lesssim h^{-1}\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0}\,$,
and using that $\kappa(\mathbb{M})\sim\mathcal{O}(1)$ (see, e.g., [EG21-II, Prop. 28.6]) we conclude that $\kappa(\mathbb{J})\sim\mathcal{O}(h^{-2})$ (and thus is independent of the value of $\beta$). This further implies
@@
from which we also conclude $\kappa(\mathbb{A}^{0})\sim\mathcal{O}\Big((1+h^{-(\beta+1)})/(1+h^{-\beta+1})\Big)$ which is bounded by $Ch^{-2}$ independently of the choice of $\beta$.
@@
$h$ $\kappa(\mathbb{A})^{-1}$ eoceoc $\kappa(\mathbb{A}^{1})^{-1}$ eoceoc $\kappa(\mathbb{A}^{0})^{-1}$ eoceoc
$0.5$ $1.52\cdot 10^{-3}$ (--) $2.03\cdot 10^{-2}$ (--) $1.28\cdot 10^{-1}$ (--)
@@
$h$ $\kappa(\mathbb{A})^{-1}$ eoceoc $\kappa(\mathbb{A}^{1})^{-1}$ eoceoc $\kappa(\mathbb{A}^{0})^{-1}$ eoceoc
$0.5$ $6.23\cdot 10^{-4}$ (--) $2.03\cdot 10^{-2}$ (--) $1.06\cdot 10^{-1}$ (--)
@@
$h$ $\kappa(\mathbb{A})^{-1}$ eoceoc $\kappa(\mathbb{A}^{1})^{-1}$ eoceoc $\kappa(\mathbb{A}^{0})^{-1}$ eoceoc
$0.5$ $9.51\cdot 10^{-5}$ (--) $2.03\cdot 10^{-2}$ (--) $6.69\cdot 10^{-2}$ (--)